Grade 12Physical Chemistry

An alkaloid contains 17.28% of nitrogen and its molecular mass is 162.How many nitrogen atoms are present in one molecule of alkaloid?

saba
9 Years agoGrade 12

3 Answers

varun
9 Years ago
HELLO,
Assume one molecule of alkaloid contains X number of nitrogen atoms
then percentage of nitrogen = 14X/162 X 100
                                           = 2
so X  = 2
SO NUMBER OF NITROGEN ATOMS IN ONE MOLECULE OF ALKALOID IS 2
THANK YOU
Nikita sharma
8 Years ago
1 mole=22.4 l = 6.022×10^23moleculesSo according to question, 22.4 l = 6.022×10^23 molecules22.4/6.02210^23 = 1molecule3.7 l = 1 molecule (ans)
RRJ
7 Years ago
Mass of Nitrogen in the compound  =17.28% * 162gm.==28g
 •.•Moles of N=28/14=2moles
 1mol alkaloid contains 2 mol N atoms
So NA molecules contain 2NA N atoms
Or 1 molecule contains 2 N atoms